The velocity of the wind over the surface of the Wing of an aeroplane is 80 ms^-1 and under the wing 60ms^-1 . If the area of the wing is 4m^2 the dynamic lift experienced by the wing is [ density of air =1.3 kg.m^-3 

Grade:11

1 Answers

Khimraj
5 years ago
Dynamic Lift = ½*pA[V22 – V12]
 = ½*1.3*4[802 – 602] = 2.6*(6400-3600) = 2.6*2800 = 7280N
Hope it clears.....................